OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation.

OpenStudy (ageta):

Substitution method is super easy. :) Since it said find the equivalent to y in the first equation... you need to isolate y! -y = -6x + 1 y = 6x - 1 Now that you have your equation, just substitute that into the second equation. 4x -3(6x - 1) = -11 4x - 18x + 3 = -11 -14x + 3 = -11 -14x = -14 now divide x = 1 YOUR answer however would be y = 6x - 1

OpenStudy (anonymous):

this is what was under the question but not part of the answers 8x - y = -9 4x - 3y = -22

OpenStudy (ageta):

yes

OpenStudy (anonymous):

4(-8x - 9) - 3y = -22 4x - 3(-8x - 9) = -22 4(8x + 9) - 3y = -22 4x - 3(8x + 9) = -22 these are the choices